# VICON: A Foundation Model for Multi-Physics Fluid Dynamics via Vision In-Context Operator Networks

### v2

@@ 4.2 Vision In-Context Operator Networks
Figure 2: Comparison of rollout errors across different datasets, using single-step and multi-step strategies with varying maximum step sizes ($s_{\text{max}}=5,6,7$).
4.3 Prompt Normalization
$\mathcal{L}^{\prime}(\mathcal{N}_{c}(\mathbf{u}))=\mathcal{N}_{q}(\mathcal{L}(\mathbf{u})),$ (8)
or equivalently:
$\mathcal{L}^{\prime}(\mathbf{v})=\mathcal{N}_{q}(\mathcal{L}(\mathcal{N}^{-1}_{c}(\mathbf{v}))).$
4.4 Datasets and Data Augmentation
The dataset used in this work contains three types of equations modeling different fluid dynamics:
  1. 1.
    PDEArena-Incomp (Gupta & Brandstetter, 2022), the incompressible Navier-Stokes (NS) equation
  2. 2.
    PDEBench-Comp-HighVis (Takamoto et al., 2022), the compressible NS equation
  3. 3.
    PDEBench-Comp-LowVis (Takamoto et al., 2022), the compressible NS equation with very low viscosity
Detailed descriptions of the datasets are provided in Appendix A.
The temporally homogeneous Markovian property (Section 4.1) enables natural data augmentation for training and flexible rollout strategy for inference (Section 4.5): by striding with larger timesteps to reduce the number of autoregressive steps in long-term prediction tasks: $\Delta t=s\Delta\tau_{i_{p}}\;|\;s=1,2,\dots,s_{\text{max}}$, where $\tau_{i_{p}}$ is the timestep size for recording in trajectory $i_{p}$. During training, for each trajectory in the dataset, we first sample a stride size $s\sim\mathcal{U}\{1,\ldots,s_{\text{max}}\}$, then form $\langle\mathbf{u}_{t},\mathbf{u}_{t+s\Delta\tau}\rangle$ pairs for the corresponding operator.
4.5 Inference with Flexible Strategies
For long-term rollout, we employ VICON in an autoregressive fashion. The ability to predict with different timestep strides enables various rollout strategies. We explore two natural approaches: single-step rollout and multi-step rollout, where the latter advances in larger strides to reduce the number of rollout steps.
For single-step rollout, we simply follow the autoregressive procedure with $s=1$. Multi-step rollout involves a more sophisticated approach: given a maximum prediction stride $s_{\text{max}}$, we first make sequential predictions with a gradually growing $s=1$ to $s_{\text{max}}$ using $\mathbf{u}_{I_{0}}$ as the question condition, obtaining $\left\{\mathbf{u}_{I_{0}+s}\right\}_{s=1}^{s_{\text{max}}}$. Using each frame in this sequence as a question condition, we then make consistent $s_{\text{max}}$-step predictions while preserving all intermediate frames. This multi-step strategy follows the approach in (Yang & Osher, 2024). Detailed examples of both rollout strategies are provided in Appendix B.3.
4.6 Evaluation Metric
We evaluate the rollout accuracy of VICON using two different strategies described in Section 4.5. The evaluation uses rescaled and absolute $L^{2}$ errors between the predicted and ground truth frames, starting from the frame $I_{0}+1$. For relative scaling coefficients, we use channel-wise scaling standard deviation $\mathbf{\sigma}$ of ground truth frames, which vary between different prompts.
(a) PDEBench-Comp-LowVis
(b) PDEBench-Comp-HighVis
(c) PDEArena-Incomp
Figure 3: Comparison of rollout errors using different patch resolutions ($8,16,32,64$).
5 Experimental Results
The detailed experimental results are presented in Table 1 and Table 8, where we report the scaled and absolute $L^{2}$ roll-out errors for the three datasets. We record the rollout error at time steps 1, 5, 10, and the last step, along with the average error across all time steps. Due to the flexibility of time steps, a single trained model adopts two rollout strategies: single-step and multistep (Section 4.5) during inference, and we report results for both strategies.
We benchmark against MPP (McCabe et al., 2023) (with the AViT-B model architecture, see details in their Table 3), the current state-of-the-art sequence-to-sequence model. The vanilla ICON, which would require processing over 114K tokens per frame (128×128×7), exceeds available GPU memory and is thus computationally infeasible for direct comparison on these datasets.
Superior Accuracy for Long-term Predictions.
We observe that VICON (both strategies) outperforms MPP at all timesteps for the PDEBench-Comp-LowVis and PDEBench-Comp-HighVis datasets. For longer rollouts, VICON shows a significant reduction in scaled $L^{2}$ error of $40\%$ and $61.6\%$ for PDEBench-Comp-LowVis and PDEBench-Comp-HighVis datasets, respectively. For the PDEArena-Incomp dataset, while VICON slightly underperforms MPP initially, it achieves better results for longer-step rollouts (the last step and the averaged error across the whole time horizon of 46 steps). In general, our method demonstrates superior long-term predictions compared to MPP.
Turbulence Kinetic Energy Analysis.
The turbulence kinetic energy (TKE), defined as $\frac{1}{2}(\overline{\tilde{u}_{x}^{2}}+\overline{\tilde{u}_{y}^{2}})$, is a critical metric to quantify the ability of a model to capture turbulent flow characteristics. Here, $\tilde{{\bm{u}}}={\bm{u}}-\overline{{\bm{u}}}$ denotes the fluctuation of the velocity from its statistical equilibrium state $\overline{{\bm{u}}}$. Within our datasets, only a subset of PDEBench-Comp-LowVis, specifically those initialized with fully developed turbulent fields (see Appendix D.5 of (Takamoto et al., 2022)), is suitable for TKE analysis. After filtering these entries, we compare the mean absolute error (MAE) of the TKE between VICON and MPP. Our approach achieves a TKE error of $0.016$, significantly outperforming the MPP’s error of $0.049$, which demonstrates superior preservation of turbulent flow characteristics. Figure 4 illustrates a representative comparison of TKE predictions between VICON and MPP.
Flexible Rollout Strategy.
To demonstrate the flexibility of our model rollout strategies (Section 4.5), we evaluate the performance of single-step and multi-step rollout strategies with varying maximum step sizes ($s_{\text{max}}=5,6,7$) in Figure 2. We observe that the model performs best with single-step rollout for PDEBench-Comp-LowVis and PDEBench-Comp-HighVis, while multi-step rollout with $s_{\text{max}}=5$ shows the best performance for PDEArena-Incomp.
@@ 5 Experimental Results
This performance difference stems from the datasets’ temporal characteristics: PDEBench-Comp-LowVis and PDEBench-Comp-HighVis use larger timesteps with fewer total frames, making accurate large-stride predictions challenging and favoring single-step approaches. In contrast, PDEArena-Incomp’s longer trajectories with smaller timesteps benefit from multi-step prediction, as reducing the number of rollout steps helps minimize error accumulation.
This flexibility in inference allows the model to adapt to different datasets and prediction tasks without retraining, making VICON more versatile for various task settings (such as climate prediction at short and long terms).
Figure 4: Comparison of turbulence kinetic energy predictions between VICON and MPP, demonstrating the improved accuracy in capturing turbulent flow features of VICON. (Top-left) Ground truth TKE field, (Left) VICON prediction and error, (Right) MPP prediction and error.
Stride Generalization.
When evaluated with unseen strides ($s_{\text{max}}=6,7$), our model maintains reasonable performance and even outperforms the single-step strategy for PDEArena-Incomp. This capability stems from ICON’s ability to implicitly learn operators via context pairs. When applied to experimental measurements, this ability to extrapolate and generalize to unseen timestep sizes provides greater tolerance for sampling rates, which can be constrained by hardware limitations.
Computational Efficiency.
@@
In Table 3, we compare the computational resources required by VICON and MPP. Our model demonstrates greater efficiency, requiring approximately one third of the inference time per frame and using 75% of the total parameters compared to MPP.
Impact of Patch Resolutions.
We conducted ablation studies on patch resolution by varying patch sizes (8, 16, 32, 64). Although smaller patches could theoretically capture finer spatial structures, they generate longer token sequences for the transformer. Due to the quadratic attention complexity, this imposes a critical trade-off: we must reduce the token dimension to stay within memory constraints. Our experiments show that reducing the patch resolution below $16\times 16$ requires lowering both the token dimension from $1024$ to $512$ and the feedforward dimension from $2048$ to $1024$ to fit within our GPU memory limits. Moreover, further reduction below the $8\times 8$ patches proved infeasible even with mixed precision training.
The results in Figure 3 demonstrate that our choice of patch size 16 achieves optimal performance in all data sets, indicating an effective balance between the length of the sequence and the representational capacity. Although performance degradation with coarser patches (32, 64) aligns with theoretical expectations, the degradation with finer patch resolution probably stems from the necessary reduction in hidden dimensions. This observation highlights the importance of carefully balancing the sequence length against hidden dimensions. The substantial computational resources required also suggest a fundamental challenge in scaling VICON to 3D applications without stronger computing infrastructure or modifications to the model.
Impact of Data Diversity on Generalization.
To evaluate the generalization capabilities of the model across datasets, we performed ablation studies by training on various combinations of datasets. The results are presented in Table 2.
The experiments reveal that the model achieves optimal performance on individual datasets when trained exclusively on them, though at the cost of poor accuracy on others. As more datasets are incorporated during training, a trade-off emerges: Performance on individual datasets slightly decreases, but the model learns to handle different simulation types simultaneously.
These results demonstrate that VICON is a generalized few-shot model for diverse fluid dynamics problems, suggesting potential extensions through additional datasets and pretrain-finetune approaches to address a much wider range of settings.
Resource and Timing Metrics Ours MPP (McCabe et al., 2023)
Training cost [GPU hrs] 58 64
Rollout time per step [ms] 8.7 25.7
Model Param Size 88M 116M
@@ 6 Conclusion and Future Works
We present VICON, a vision-transformer-based in-context operator network that efficiently handles two-dimensional fluid dynamics problems. By processing dense physical fields through patch-wise operations, VICON overcomes the computational limitations of naive ICON approaches while maintaining their flexibility in multi-physics operator learning.
In extensive experiments, we have demonstrated that VICON achieves superior performance compared to existing MPP, especially in long-term predictions. VICON also offers flexible rollout strategies with varying time-step strides, allowing users to adapt predictions based on specific scenario requirements.
Several promising directions remain for future investigation. Our method currently requires full domain input. 1. Incorporating physical priors, such as translational invariance, could enable domain decomposition for training on smaller subdomains, effectively decoupling computational complexity from domain size. 2. Extending VICON to handle irregular domains and different data structures such as graphs or point clouds would broaden its applicability to more realistic scenarios. 3. Our current approach of joining channels across datasets has limited scalability. Drawing inspiration from machine translation’s subword tokenization, developing adaptive physical field representations could offer a more scalable cross-dataset solution. 4. Most critically, extending to 3D applications faces fundamental challenges as token sequence length grows cubically with spatial dimensions, while memory constraints limit the hidden dimensions and representational capacity. Novel architectures combining in-context learning with efficient data compression techniques may be necessary to address this limitation.
@@ Acknowledgment
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 PDEArena-Incomp Dataset
The incompressible Navier-Stokes dataset comes from PDEArena (Gupta & Brandstetter, 2022). The data are generated from the equation
@@ A.2 PDEBench-Comp-HighVis and PDEBench-Comp-LowVis Datasets
The PDEBench-Comp-HighVis and PDEBench-Comp-LowVis datasets come from PDEBench Compressible Navier-Stokes dataset (Takamoto et al., 2022). The data are generated from the equation
@@ B.2 Training Details
We implement our method in PyTorch (Paszke et al., 2019) and utilize data parallel training (Li et al., 2020) across two NVIDIA RTX 4090 GPUs. We employ the AdamW optimizer with a cosine learning rate schedule that includes a linear warmup phase. We apply gradient clipping with a maximum norm of 1.0 to ensure training stability. All optimization parameters are detailed in Table 5.
@@
B.3 Rollout Strategy Example
We demonstrate two rollout strategies in Tables 6 and 7, showing single-step and multi-step strategies, respectively. In both cases, the initial frames span from time step 0 to 9, and we aim to predict the trajectory up to time step 20.
@@ B.3 Rollout Strategy Example
Rollout index Example CONDs Example QOIs Question COND Predict QOI
1 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 9 10
2 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 10 11
3 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 11 12
4 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 12 13
5 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 13 14
6 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 14 15
7 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 15 16
8 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 16 17
9 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 17 18
10 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 18 19
11 [0,1,2,3,4,5,6,7,8] [1,2,3,4,5,6,7,8,9] 19 20
@@
As shown in Table 7, the multi-step strategy initially uses smaller strides to build sufficient examples, then employs maximum strides for later rollouts, as detailed in Section 4.5. We note that repeated in-context examples appear in Table 7, which is common when the maximum stride is large and the initial frames cannot form enough examples. While the number of in-context examples in VICON is flexible and our model can accommodate fewer examples than the designed length, our preliminary experiments indicate that the model performs better with more in-context examples, even when some examples are repeated.

### v3

@@ 5 Experimental Results
Figure 3: Comparison of rollout errors (scaled by std) across different datasets and models. We show errors for two VICON rollout strategies: single step rollout and flexible rollout strategy. For flexible step, step 3 works optimally for the PDEArena-Incomp dataset, while step 1 works best for PDEBench-Comp-LowVis and PDEBench-Comp-HighVis.
5.1 Superior Performance on Long-term Rollout Predictions
As demonstrated in Figure 2(a), Table 1, and Figure 3, VICON consistently outperforms baseline methods on long-horizon predictions across all benchmarks—with the exception of DPOT on the PDEArena-Incomp dataset, where our performance is comparable. Overall, VICON achieves an average reduction in relative $L^{2}$ RMSE at the final timestep of 37.9% compared to DPOT and 44.7% compared to MPP.
Notably, DPOT exhibits exceptionally poor performance on the PDEBench-Comp-HighVis dataset, with an 821.9% error compared to our 30.06%. Our visualization of failure cases reveals that DPOT struggles with trajectories with small pressure values compared to the dataset average. This may stems from DPOT’s lack of prompt normalization, as compressible flow’s pressure channel exhibit large magnitude variations.
For the PDEArena-Incomp dataset, while VICON slightly underperforms DPOT and MPP initially, it quickly surpasses MPP and achieves comparable performance to DPOT for longer-step rollouts. This demonstrates VICON’s robustness in long-term predictions. Despite DPOT’s marginally better performance here, its poor performance on the PDEBench-Comp-HighVis dataset limits its applicability in multi-physics settings.
More detailed results are provided in Table 8 and Table 9 in Appendix D.
5.2 Flexible Rollout Strategies
As demonstrated in Figure 2(b) and Figure 5, VICON can select appropriate timestep strides based on each dataset’s characteristics. While PDEBench-Comp-LowVis and PDEBench-Comp-HighVis perform best with single-step rollout, PDEArena-Incomp achieves optimal results with a stride of 3 using our flexible-step rollout strategy.
@@ 5.2 Flexible Rollout Strategies
This dataset-dependent performance is expected: PDEBench-Comp-LowVis and PDEBench-Comp-HighVis record data with a larger timestep size and fewer total frames, making multi-stride predictions challenging. Conversely, PDEArena-Incomp records data with smaller timesteps and benefits from multi-stride prediction, as this approach reduces the rollout steps and error accumulation. As a result, VICON achieves an 11.4% error reduction in PDEArena-Incomp by selecting a balanced stride ($s_{\text{max}}=3$).
When evaluated with unseen strides ($s_{\text{max}}=7$), VICON maintains comparable performance to single-step strategies for PDEArena-Incomp, indicating that it extracts underlying operators through context pairs rather than memorizing dynamics. This generalization capability provides tolerance when deployed to real-world settings where device sampling rates often differ from a fixed training set—an advantage we further demonstrate in the next section.
5.3 Robustness to Imperfect Temporal Measurements
Real-world experimental measurements frequently suffer from imperfections—sampling rates may differ from training set, and frames can be missing due to device errors. Sequence-to-sequence models like MPP and DPOT require either retraining with new data or interpolation that introduces noise. In contrast, VICON elegantly handles such imperfections by forming context pairs from available frames only (Figure 1(c)). The algorithm for generating pairs with imperfect measurements appears in Section C.2.
We evaluate VICON against baselines on all benchmarks under two scenarios: (1) half sampling rate (dropping every other frame) and (2) random frame dropping (removing 1-3 frames per sequence). Results for the PDEBench-Comp-LowVis dataset in Figure 2(c) (with complete results in Figure 8) show that VICON experiences only 24.41% relative performance degradation compared to 71.37%-74.49% in baselines—demonstrating remarkable robustness to measurement imperfections.
5.4 Turbulence Kinetic Energy Analysis
Turbulence kinetic energy (TKE), defined as $\frac{1}{2}(\overline{\tilde{u}_{x}^{2}}+\overline{\tilde{u}_{y}^{2}})$, is a critical metric for quantifying a model’s ability to capture turbulent flow characteristics. Here, $\tilde{{\bm{u}}}={\bm{u}}-\overline{{\bm{u}}}$ denotes the fluctuation of velocity from its statistical equilibrium state $\overline{{\bm{u}}}$.
Within our datasets, only a subset of PDEBench-Comp-LowVis, specifically those initialized with fully developed turbulent fields (see Appendix D.5 of [49]), is suitable for TKE analysis. After filtering these entries, we compare the mean absolute error (MAE) of the TKE between VICON, DPOT, and MPP. Our approach achieves a TKE error of $0.016$, significantly outperforming MPP’s error of $0.049$, while achieving performance comparable to DPOT’s error of $0.012$. Visualizations of TKE errors are presented in Figure 6 in Appendix D.
5.5 Benefits of Multi-Physics Joint Training
We empirically examine whether VICON benefits from a multi-physics training by comparing two strategies: (1) training a single model jointly on all three datasets versus (2) training separate models specialized for individual datasets. For fair comparison, we maintained identical batch sizes across all trainings while adjusting the steps for separate training to be slightly more than one-third of the joint training duration, ensuring comparable total computational costs between approaches.
As shown in Figure 7, joint training significantly outperforms separate training across all three datasets. We attribute this performance gain to the underlying physical similarities across these fluid dynamics problems, where exposure to diverse yet related flow patterns enhances the model’s generalization ability.
5.6 Computational Efficiency
@@
In Figure 2(d) and Table 11, we compare the computational resources required by VICON and baselines. Our model demonstrates superior efficiency across multiple metrics. Compared to MPP, VICON requires approximately one-third of the inference time per frame while using 75% of the total parameters. VICON also outperforms DPOT, requiring 28% fewer parameters and 28% less inference time.
@@ 6 Conclusion and Future Works
We present VICON, a vision-transformer-based in-context operator network that efficiently processes dense physical fields through patch-wise operations. VICON overcomes the computational burden of original in-context operator networks in higher spatial dimensions while preserving the flexibility to extract multiphysics dynamics from few-shot contexts.
Our comprehensive experiments demonstrate that VICON achieves superior performance in long-term predictions, reducing the averaged last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times. The model supports flexible rollout strategies with varying timestep strides, enabling natural application to imperfect real-world measurements where sampling frequencies differ or frames are randomly dropped—scenarios where VICON experiences only 24.41% relative performance degradation compared to 71.37%-74.49% degradation in baseline models.
Despite these advances, several challenges remain for future investigation. We emperically showed in Section 5.5 the benefits of multi-physics training, yet scaling to larger and more diverse datasets presents practical challenges. Specifically, generating high-fidelity physics simulation data across multiple domains requires both substantial computational resources and specialized domain expertise. Furthermore, the current approach does not yet extend to 3D applications, as token sequence length would grow cubically, exceeding our computational budget. The channel-union approach is also limited when incorporating domains beyond fluid dynamics with fundamentally different state variables. Finally, adapting VICON to handle irregular domains such as graphs or meshes would broaden its applications to areas like solid mechanics and molecular dynamics. Addressing these challenges will advance the promising paradigm of in-context learning for a broader range of physical systems.
@@ A.1 PDEArena-Incomp Dataset
The incompressible Navier-Stokes dataset comes from PDEArena [1]. The data are generated from the equation
@@ A.2 PDEBench-Comp-HighVis and PDEBench-Comp-LowVis Datasets
The PDEBench-Comp-HighVis and PDEBench-Comp-LowVis datasets come from PDEBench Compressible Navier-Stokes dataset [49]. The data are generated from the equation
@@ B.2 Training Details
We implement our method in PyTorch [50] and utilize data parallel training [51] across two NVIDIA RTX 4090 GPUs. We employ the AdamW optimizer with a cosine learning rate schedule that includes a linear warmup phase. We apply gradient clipping with a maximum norm of 1.0 to ensure training stability. All optimization parameters are detailed in Table 3.
@@
B.3 MPP Details
For a fair comparison, we retrained MPP [26] using our training dataset using the same model configurations and optimizer hyperparameters (batch size is set to be the maximum possible on our device). We evaluate MPP using the same testing setup and metric.
B.4 DPOT Details
For a fair comparison, we retrain DPOT [27] using our training dataset using the same model configurations and optimizer hyperparameters (batch size is set to be the maximum possible on our device). We evaluate DPOT using the same testing setup and metric.
Appendix C Algorithms and Examples of Strategy Generation
C.1 Strategy with Full Temporal Sequence
Algorithm for single-step strategy.
Algorithm 1 presents our approach for generating single-step rollout strategies when all temporal frames are available. This strategy maintains a fixed stride of 1 between consecutive frames, providing stable but potentially error-accumulating predictions.
Input: D: Number of demonstrations to use
R: Number of ground truth reference steps
T: Total steps in sequence
Output: S: Strategy list of example pairs and question pairs
1
2S $\leftarrow$ [ ]
3
Ei $\leftarrow$ range(0, D)
  /* example input indices: 0,1,...,D-1 */
Eo $\leftarrow$ range(1, D+1)
  /* example output indices: 1,2,...,D */
4
5for i from R to T-1 do
       Qi $\leftarrow$ i - 1
        /* question input is previous frame */
       Qo $\leftarrow$ i
        /* question output is current frame */
6       S.append((Ei, Eo), Qi, Qo)
7
8 end for
9
10return S
Algorithm 1 GenerateSingleStepStrategy
Algorithm for flexible-step strategy.
Algorithm 2 describes our approach for generating flexible-step strategies with variable strides up to a maximum value. This approach reduces the total number of rollout steps required, potentially mitigating error accumulation for long sequences.
Input: D: Number of demonstrations to use
R: Number of ground truth reference steps
M: Maximum stride between pairs
T: Total steps in sequence
Output: S: Strategy list of example pairs and question pairs
1
/* Require R >= M + 1 to ensure examples exist for each stride */
2 S $\leftarrow$ [ ]
[…]
  /* example condition indices by stride */
[…]
  /* example query indices by stride */
3
/* Prepare example pairs for each stride */
4 for s from 1 to M do
       Nd $\leftarrow$ min(D, R - s)
        /* available examples for this stride */
       Eqo $\leftarrow$ range(R - Nd, R)
        /* output indices for examples */
5
      /* If we need more examples than available, repeat them */
[…]
        /* ceiling division */
       Eqo $\leftarrow$ repeat(Eqo, reps)[0:D]
        /* repeat and truncate to D */
       Eqo.sort()
        /* ensure increasing order */
6
      Ec[s] $\leftarrow$ Eqo - s
        /* input indices are s steps before outputs */
       Eq[s] $\leftarrow$ Eqo
        /* store output indices */
7
8 end for
9
/* Generate rollout strategy */
10 for i from R to T-1 do
       dist $\leftarrow$ i - R + 1
        /* distance from last reference frame */
       s $\leftarrow$ min(dist, M)
        /* select appropriate stride */
11
      Qi $\leftarrow$ i - s
        /* question input index */
       Qo $\leftarrow$ i
        /* question output index */
12
13      S.append((Ec[s], Eq[s]), Qi, Qo)
14
15 end for
16
17return S
Algorithm 2 GenerateFlexibleStepStrategy
Rollout Strategy Example.
We demonstrate two rollout strategies in Tables 4 and 5, showing single-step and flexible-step strategies, respectively. In both cases, the initial frames span from time step 0 to 9, and we aim to predict the trajectory up to time step 20.
@@ 2 S $\leftarrow$ [ ]
Rollout index Examples (COND, QOI) Question COND Predict QOI
1 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 9 10
2 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 10 11
3 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 11 12
4 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 12 13
5 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 13 14
6 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 14 15
7 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 15 16
8 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 16 17
9 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 17 18
10 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 18 19
11 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 19 20
@@
As shown in Table 5, the flexible-step strategy initially uses smaller strides to build sufficient examples, then employs maximum strides for later rollouts, as detailed in Section 4.5. We note that repeated in-context examples appear in Table 5, which is common when the maximum stride is large and the initial frames cannot form enough examples. While the number of in-context examples in VICON is flexible and our model can accommodate fewer examples than the designed length, our preliminary experiments indicate that the model performs better with more in-context examples, even when some examples are repeated.
Table 5: Flexible-step Rollout Strategy Example ($s_{\text{max}}=5$)
Rollout index Examples (COND, QOI) Question COND Predict QOI
1 (0,1) (1,2) (2,3) (3,4) (4,5) (5,6) (6,7) (7,8) (8,9) 9 10
2 (0,2) (0,2) (1,3) (2,4) (3,5) (4,6) (5,7) (6,8) (7,9) 9 11
3 (0,3) (0,3) (1,4) (1,4) (2,5) (3,6) (4,7) (5,8) (6,9) 9 12
4 (0,4) (0,4) (1,5) (1,5) (2,6) (2,6) (3,7) (4,8) (5,9) 9 13
5 (0,5) (0,5) (1,6) (1,6) (2,7) (2,7) (3,8) (3,8) (4,9) 9 14
6 (0,5) (0,5) (1,6) (1,6) (2,7) (2,7) (3,8) (3,8) (4,9) 10 15
7 (0,5) (0,5) (1,6) (1,6) (2,7) (2,7) (3,8) (3,8) (4,9) 11 16
8 (0,5) (0,5) (1,6) (1,6) (2,7) (2,7) (3,8) (3,8) (4,9) 12 17
9 (0,5) (0,5) (1,6) (1,6) (2,7) (2,7) (3,8) (3,8) (4,9) 13 18
10 (0,5) (0,5) (1,6) (1,6) (2,7) (2,7) (3,8) (3,8) (4,9) 14 19
11 (0,5) (0,5) (1,6) (1,6) (2,7) (2,7) (3,8) (3,8) (4,9) 15 20
C.2 Strategy with Imperfect Temporal Sequence
When facing imperfect temporal sampling where certain frames are missing, we can still form valid demonstration pairs from the available frames. Algorithm 3 presents this adaptive pair selection process, which serves as a fundamental building block for all strategy generation algorithms in scenarios with irregular temporal data.
Input: D: Number of demonstrations needed
dt: Desired time stride between pairs
Fa: List of indices of available frames
Output: P: List of input-output index pairs with the specified stride
1
2Fa.sort()
P $\leftarrow$ [ ]
  /* Initialize empty pairs list */
3
/* Find all pairs with stride dt */
4 for i from 0 to len(Fa) - 1 do
5       for j from i + 1 to len(Fa) - 1 do
6             if Fa[j] - Fa[i] == dt then
7                   P.append((Fa[i], Fa[j]))
8
9             end if
10
11       end for
12
13 end for
14
15if P is empty then
16       return [ ]
17
18 end if
19
/* If we don’t have enough unique pairs, repeat them */
20 if len(P) < D then
[…]
        /* ceiling division */
       RP $\leftarrow$ repeat(P, reps)[0:D]
        /* repeat pairs and truncate to D */
21       P $\leftarrow$ RP
22
23 end if
24
25return P
Algorithm 3 GetAvailablePairs
Algorithm for single-step strategy with drops.
Algorithm 4 extends the single-step strategy to handle scenarios where frames are missing from the input sequence, adaptively forming strategies from available frames while maintaining the fixed stride constraint.
Input: D: Number of demonstrations to use
S: Fixed stride between pairs
T: Total steps in sequence
Fa: List of indices of available frames
Output: St: Strategy list of (E, Qi, Qo) tuples
1
2St $\leftarrow$ [ ]
3 Fa.sort()
Fs $\leftarrow$ Fa[-1]
  /* the starting frame for rollout */
4
P $\leftarrow$ GetAvailablePairs(D, S, Fa)
  /* pairs with fixed stride */
5 if P is empty then
       return [ ]
        /* No available pairs with the specified stride */
6
7 end if
8
Fc $\leftarrow$ Fa.copy()
  /* accumulated frames (current + predicted) */
9
10for i from (Fs + 1) to (T - 1) do
       Ci $\leftarrow$ i - S
        /* potential condition index */
11
12      if Ci $\not\in$ Fc then
             continue
              /* Cannot predict frame i with stride S */
13
14       end if
15
16      Pdt $\leftarrow$ P[s]
        /* The example pairs for predicting frame i */
       Qi $\leftarrow$ i - s
        /* question input index */
       Qo $\leftarrow$ i
        /* question output index */
17
18      S.append(Pdt, Qi, Qo)
       Fc.append(Qo)
        /* The predicted frame can be used for future steps */
19
20 end for
21
22return St
Algorithm 4 GenerateSingleStepStrategyWithDrops
Algorithm for flexible-step strategy.
Algorithm 5 presents our solution for generating flexible-step rollout strategies when frames are missing, dynamically selecting appropriate strides based on available frames and previous predictions.
Input: D: Number of demonstrations to use
M: Maximum stride between pairs for rollout
T: Total steps in the sequence
Fa: List of indices of available frames
Output: S: Strategy list of (E, Qi, Qo) tuples, where E are the list of example pair indices (Ei, Eo) and Qi, Qo are the question and output indices
1 S $\leftarrow$ [ ]
2 Fa.sort()
Fs $\leftarrow$ Fa[-1]
  /* the starting frame for rollout */
3
[…]
  /* dictionary of available pairs */
4 for dt from 1 to M do
       Pdt $\leftarrow$ GetAvailablePairs(D, dt, Fa)
        /* pairs with stride dt in Fa */
5       if Pdt is not empty then
6             P[dt] $\leftarrow$ Pdt
7
8       end if
9
10 end for
Ms $\leftarrow$ max(P.keys())
  /* maximum available stride */
Fc $\leftarrow$ Fa.copy()
  /* accumulated frames (current + predicted) */
11
12for i from (Fs + 1) to (T - 1) do
13       dt $\leftarrow$ i - Fs
       Mt $\leftarrow$ min(dt, M, Ms)
        /* maximum stride for current step */
14
15      found $\leftarrow$ False
16       for s in P.keys().sorted(reverse=True) where s $\leq$ Mt do
             Ci $\leftarrow$ i - s
              /* potential condition index */
17             if Ci $\in$ Fc then
                   /* Found a starting index to obtain frame i */
18                   found $\leftarrow$ True
19                   break
20
21             end if
22
23       end for
24
25      if not found then
             /* We cannot predict frame i */
26             continue
27
28       end if
29
      /* The example pairs for predicting frame i */
30       Pdt $\leftarrow$ P[s]
       Qi $\leftarrow$ i - s
        /* question input index */
       Qo $\leftarrow$ i
        /* question output index */
31
32      S.append(Pdt, Qi, Qo)
33       Fc.append(Qo)
        /* The Qo frame is obtained and can be used for future rollouts */
34
35 end for
36
return S
Algorithm 5 GenerateFlexibleStepStrategyWithDrops
Rollout Strategy Example with Missing Frames.
Tables 6 and 7 illustrate our adaptive strategies when frames 2, 5, and 9 are missing from the initial sequence (with all other settings identical to those in Section C.1). The tables demonstrate single-step and flexible-step (max stride: 3) approaches, respectively, highlighting how our algorithm dynamically selects appropriate example pairs to maintain prediction capability despite missing some temporal samplings.